OpenStudy (mythorius):

Using the technique in the model above, find the missing segments in this 30°-60°-90° right triangle. AD = 6 DB = 2 AC =

OpenStudy (mythorius):

@KamiBug

OpenStudy (welshfella):

Triangle ACD is 30-60-90 triangle so side AD = 6 = sqrt3 x giving x = 6 / sqrt3 so CD = 6 / sqrt3 and AC = 2*6 / sqrt3

OpenStudy (welshfella):

rationalising the denominator AC = 12 sqrt3 / 3 = 4 sqrt3

OpenStudy (welshfella):

the special triangle 30-60-90 is worth remembering The sides are always in this ratio